# MOOSE-Chem: Large Language Models for Rediscovering Unseen Chemistry Scientific Hypotheses

### v5

@@ Abstract
Scientific discovery plays a pivotal role in advancing human society, and recent progress in large language models (LLMs) suggests their potential to accelerate this process. However, it remains unclear whether LLMs can autonomously generate novel and valid hypotheses in chemistry. In this work, we investigate whether LLMs can discover high-quality chemistry hypotheses given only a research background—comprising a question and/or a survey—without restriction on the domain of the question. We begin with the observation that hypothesis discovery is a seemingly intractable task. To address this, we propose a formal mathematical decomposition grounded in a fundamental assumption: that most chemistry hypotheses can be composed from a research background and a set of inspirations. This decomposition leads to three practical subtasks—retrieving inspirations, composing hypotheses with inspirations, and ranking hypotheses—which together constitute a sufficient set of subtasks for the overall scientific discovery task. We further develop an agentic LLM framework, MOOSE-Chem, that is a direct implementation of this mathematical decomposition. To evaluate this framework, we construct a benchmark of 51 high-impact chemistry papers published and online after January 2024, each manually annotated by PhD chemists with background, inspirations, and hypothesis. The framework is able to rediscover many hypotheses with high similarity to the groundtruth, successfully capturing the core innovations—while ensuring no data contamination since it uses an LLM with knowledge cutoff date prior to 2024. Finally, based on LLM’s surprisingly high accuracy on inspiration retrieval, a task with inherently out-of-distribution nature, we propose a bold assumption: that LLMs may already encode latent scientific knowledge associations not yet recognized by humans. Contribution during internship at Shanghai Artificial Intelligence Laboratory. Corresponding author.
@@ 1 Introduction
We design experiments with the benchmark to test the three fundamental questions and find that LLMs are highly capable. We also test MOOSE-Chem with the benchmark, mimicking the setting to run it in the wild by only giving a background and a corpus of up to 3000 chemistry papers to select inspiration. Even in this challenging setting, MOOSE-Chem can still rediscover many hypotheses with very high similarity with the ground truth ones, covering the main innovations.
@@
  • We provide the first mathematical derivation on how to decompose the seemingly impossible-to-solve question $P(\text{hypothesis}|\text{research background})$ into many executable and practical smaller steps. This decomposition make $P(\text{hypothesis}|\text{research background})$ possible to be practical.
@@ 3 Benchmark Construction
To achieve the goals, we split each collected paper into the following components: $<$background question, background question (strict), background survey, background survey (strict), one to three inspiration paper titles and their reason to serve as an inspiration, research hypothesis, experiments, reasoning process, summarization of inspirations$>$. Every component is described by text.
@@ 4.1 Fundamental Assumption and Following Decomposition
$P(h\mid b)\approx\prod_{j=1}^{k}P(i_{j}\mid b,h_{j-1},I)\cdot P(h_{j}\mid b,i_{j},h_{j-1}),\text{ where }h_{0}=\emptyset$ (2)
Here, $I$ denotes the full (chemistry) literature, representing the full inspiration space to search for every single $i$. The full proof along with detailed analyses is shown in § A.1, which is the core of this paper (and therefore highly recommend to take a read).
Equation 2 is meaningful in that by decomposing $P(h\mid b)$ into more practical and executable smaller questions, the seemingly impossible-to-solve $P(h\mid b)$ itself becomes practical. We analyze how $P(i_{j}\mid b,h_{j-1},I)$ and $P(h_{j}\mid b,i_{j},h_{j-1})$ are practical and executable by LLMs in § 5.1 and § 5.2 correspondingly.
@@
Supported by Equation 2 and Equation 3, as a result, to model $P(h\mid b)$, the only three components we need to model are $P(i_{j}\mid b,h_{j-1},I)$, $P(h_{j}\mid b,i_{j},h_{j-1})$, and $R(h)$. The implementation details of the three components are illustrated in the remaining subsections in § 4. Analyses of LLM’s ability on the three components are provided in § 5.
@@ 4.2.1 The General Picture
Our methodology is developed based on the fundamental assumption discussed in § 4.1. Specifically, we use LLMs to perform $P(i_{j}\mid b,h_{j-1},I)$, $P(h_{j}\mid b,i_{j},h_{j-1})$, and $R(h)$, and organize them into a multi-agent LLM-based framework. The input to the framework is only a background question and/or background survey, together with a (large) chemistry literature corpus to search for inspiration. The output of the framework is a list of ranked research hypothesis.
@@ 4.2.2 Design Details of $P(i_{j}\mid b,h_{j-1},I)$ And Its Motivation
4.2.3 Design Details of $P(h_{j}\mid b,i_{j},h_{j-1})$ And Its Motivation
@@ 5 Investigation on Fundamental Questions
$P(h\mid b)$ can be understood as the task to discover high-quality chemistry research hypothesis, given only a background question and/or background survey. Our central question to investigate is how well LLMs can perform $P(h\mid b)$. Supported by Equation 2 and 3, we break up this main question into three smaller questions: how well can LLMs perform (1) $P(i_{j}\mid b,h_{j-1},I)$, (2) $P(h_{j}\mid b,i_{j},h_{j-1})$, and (3) $R(h)$? All experiments are performed by GPT-4o (its training data is up to October 2023).
@@ 5.1 How Well can LLMs perform $P(i_{j}\mid b,h_{j-1},I)$?
5.2 How Well can LLMs perform $P(h_{j}\mid b,i_{j},h_{j-1})$?
5 points Generated hypothesis covers three key points (or covers all the key points) and leverage them similarly as in the groundtruth hypothesis; Extra key points do not have apparent flaws.
4 points Generated hypothesis covers three key points (or covers all the key points) and leverage them similarly as in the groundtruth hypothesis; Extra key points have apparent flaws.
3 points Generated hypothesis covers two key points and leverage them similarly as in the groundtruth hypothesis, but does not cover more or all key points
2 points Generated hypothesis covers one key point and leverage it similarly as in the groundtruth hypothesis, but does not cover more or all key points
1 point Generated hypothesis covers at least one key point, but is used differently as in the groundtruth hypothesis
0 point Generated hypothesis does not cover any key point
@@ 6.1 Baselines
is a hypothesis discovery framework for the NLP and biochemical domain. It relies on semantic and citation neighbors to retrieve information to assist the background. As a result, the retrieved information could be very related to the background that might not be able to serve as an inspiration. To make the generated hypothesis more novel, it adopts self-refine to focus on improving the novelty aspect of the generated hypothesis. Here, we implement SciMON with LLM-based inspiration retrieval, the same as MOOSE-Chem. Table 3 shows that the recall rate of LLM-based retrieval is 83.7%.
@@ 6.2 Results
Table 11: MOOSE-Chem runs with $|I|$=300, mimicking the copilot setting. This table shows the statistics of the top Matched Score across the benchmark. The evaluation is done by experts.
@@ 7 Conclusion
We investigated this central question: “Can LLMs automatically discover novel and valid chemistry (including material science) research hypotheses (even those which deserve a publication in Nature, Science, or a similar level) given only a chemistry research background (consisting of a research question and/or a background survey), without limitation on the domain of the research question?”. We proposed a fundamental assumption to break up this seemingly impossible-to-solve central question into three smaller, more practical, and executable fundamental questions. Then, we investigated LLM’s ability on each of them.
To this end, we constructed a benchmark consisting of chemistry and material science papers published and only be public in 2024. We also developed an LLM-based multi-agent framework consisting of three stages reflecting the three smaller fundamental questions. Experiments showed that the framework (runs in a copilot in-the-wild setting, with LLMs with training data up to October 2023) can rediscover many hypotheses with very high similarity with the ground-truth ones, covering the main innovations.
@@ Appendix A Appendix
A.1 Proof of the Rigorous Decomposition of $P(h\mid b)$ Based on the Fundamental Assumption
@@
This assumption is reminiscent of Swanson Linking (Swanson, 1986) in the domain of literature-based discovery (LBD), also known as the “ABC model”, where two concepts A and C are hypothesized as linked if they both co-occur with some intermediate concept B in papers. Our assumption differs in that: (1) for a chemistry hypothesis published in a good venue, usually more than one inspirations are needed; (2) background and inspiration are not necessarily linked by a path of intermediate papers; (3) our assumption is applied to a majority of existing published chemistry hypotheses, while LBD has been considered to only focus on a very specific, narrow type of hypothesis (Wang et al., 2024b). It might indicate that a similar proportion of future chemistry hypotheses can also be resulted from linkages of existing literature.
@@
Here, $k\in\mathbb{Z}$ represents the number of inspirations needed for a particular $h$. Typically in chemistry, $k\in[1,3]$.
@@
$P(h\mid b)$ $=P(h,i_{1},\ldots,i_{k}\mid b)$ (5)
=$=\begin{cases}\frac{P(h,b,i_{1})}{P(b,i_{1})}\cdot\frac{P(b,i_{1})\cdot P(I)}{P(b)\cdot P(I)}&\text{if }k=1\\ \frac{P(h,b,i_{1},\ldots,i_{k})}{P(b,i_{1},\ldots,i_{k})}\cdot\frac{P(b,i_{1},\ldots,i_{k})\cdot P(I)}{P(b,i_{1},\ldots,i_{k-1})\cdot P(I)}\cdot\ldots\cdot\frac{P(b,i_{1})\cdot P(I)}{P(b)\cdot P(I)}&\text{if }k>1\\ \end{cases}$ (6)
=$=\begin{cases}P(h\mid b,i_{1})\cdot P(i_{1}\mid b,I)&\text{if }k=1\\ P(h\mid b,i_{1},\ldots,i_{k})\cdot\prod_{j=2}^{k}P(i_{j}\mid b,i_{1},\ldots,i_{j-1},I)\cdot P(i_{1}\mid b,I)&\text{if }k>1\end{cases}$ (7)
Here, $P(h,i_{1},\ldots,i_{k}\mid b)$ can be interpreted as the joint distribution over $h$ and its associated inspirations conditioned on background knowledge $b$. Equation 5 holds because Equation 4 asserts that $\{i_{1},\ldots,i_{k}\}$ constitutes a necessary and sufficient set of additional knowledge required to compose $h$. $I$ denotes the full inspiration space—that is, the set of all possible knowledge pieces that could serve as an inspiration for generating a new hypothesis. This space includes not only all existing chemistry knowledge but also potentially relevant knowledge from other disciplines that could be leveraged in formulating novel chemistry hypotheses. However, computing over the full space $I$ is computationally infeasible. To make the problem tractable, we approximate $I$ with a large but manageable subset $\hat{I}$, consisting of approximately 3,000 top cited chemistry papers from the existing chemistry literature.
@@
$P(h_{k}\mid b,i_{1},\ldots,i_{k})$ $\approx P(h_{k}\mid b,f(b,i_{1},\ldots,i_{k-1}),i_{k})$ if $\text{if }k>1$ (8)
$=P(h_{k}\mid b,h_{k-1},i_{k})$ if $\text{if }k>1$ (9)
@@
$P(i_{k+1}\mid b,i_{1},\ldots,i_{k},I)$ $\approx P(i_{k+1}\mid b,f(b,i_{1},\ldots,i_{k}),I)$ if $\text{if }k>1$ (10)
$=P(i_{k+1}\mid b,h_{k},I)$ if $\text{if }k>1$ (11)
As a result, to achieve the final $h_{k}$, we need to obtain $\{h_{1},\ldots,h_{k-1}\}$ first (if $k>1$). In addition, seeing $h$ as a “state”, and $i$ as an “action”, obtaining $h$ and $i$ through $P(h_{k}\mid b,h_{k-1},i_{k})$ and $P(i_{k+1}\mid b,h_{k},I)$ correspondingly indicates a Markov property: (1) a new $h$ only depends on $b$, its previous $h$, and the current $i$; and (2) an $i$ only depends on $b$, $I$, and the previous $h$.
For brevity, we visualize this as:
@@
where each transition is still conditioned on the background knowledge $b$, though $b$ is omitted from the notation to emphasize the Markov structure of the progression.
Building on this visualization, we formalize the formation of a hypothesis $h$ (specifically, $h=h_{k}$) as a constructive process that sequentially integrates a set of inspirations $\{i_{1},\ldots,i_{k}\}$ into intermediate hypothesis states $\{h_{1},\ldots,h_{k}\}$.
Formally, the conditional probability $P(h\mid b)$ can be expressed as a marginal over all valid sequences of inspirations that can generate $h$:
@@
where $\Pi_{k}$ denotes the set of all permutations of $\{1,\ldots,k\}$ applied to the inspirations $\{i_{1},\ldots,i_{k}\}$ such that the resulting composition yields the final hypothesis $h_{k}=h$, under the assumption in Equation 4, which specifies that $h$ is fully determined by $b$ and $\{i_{1},\ldots,i_{k}\}$. In this context, $h_{j}$ represents the intermediate hypothesis state obtained after integrating $i_{\pi(j)}$ at step $j$.
The degree to which the order of inspirations $\{i_{1},\ldots,i_{k}\}$ affects hypothesis formulation can vary across disciplines. In empirical sciences such as chemistry, the contributions of individual inspirations are largely interchangeable, and their order of integration has limited impact on the final hypothesis, resulting in a large $|\Pi_{k}|$. Conversely, in disciplines like mathematics, where hypotheses (e.g., theorems) often require constructing a specific sequence of lemmas and prior results, the ordering of inspirations is more constrained and may follow a near-deterministic path.
Considering this variance depending on the disciplines, to make this problem tractable and easier to understand, we adopt a fixed constructive path assumption, i.e., $|\Pi_{k}|=1$, selecting a representative order of inspirations $\{i_{1},\ldots,i_{k}\}$.
Under this inspiration-fixed-order assumption, where hypothesis $h$ is constructed through a specific sequence of inspirations $\{i_{1},\ldots,i_{k}\}$, we have:
@@
Therefore, if with this inspiration-fixed-order assumption for simplicity, if $k>1$,
@@
$=\frac{P(b,i_{1},I)}{P(b,I)}\cdot\frac{P(b,i_{1},h_{1})}{P(b,i_{1})}\cdot\ldots\cdot\frac{P(b,i_{k},h_{k-1},I)}{P(b,h_{k-1},I)}\cdot\frac{P(b,i_{k},h_{k-1},h_{k})}{P(b,i_{k},h_{k-1})}$ (17)
$=P(i_{1}\mid b,I)\cdot P(h_{1}\mid b,i_{1})\cdot\prod_{j=1}^{k-1}P(i_{j+1}\mid b,h_{j},I)\cdot P(h_{j+1}\mid b,i_{j+1},h_{j})$ (18)
$=\prod_{j=1}^{k}P(i_{j}\mid b,h_{j-1},I)\cdot P(h_{j}\mid b,i_{j},h_{j-1}),\text{ where }h_{0}=\emptyset$ (19)
Although starting from $k>1$, Derivation 19 covers the situation when $k=1$ in Equation 7.
Therefore, in sum, we break up the seemingly impossible question $P(h\mid b)$ into many practical and executable smaller questions as:
$P(h\mid b)\approx\prod_{j=1}^{k}P(i_{j}\mid b,h_{j-1},I)\cdot P(h_{j}\mid b,i_{j},h_{j-1}),\text{ where }h_{0}=\emptyset\text{ and }k>=1$ (20)
Of course, without the inspiration-fixed-order assumption, a more complete derivation of $P(h\mid b)$ involves marginalizing over all valid permutations in $\Pi_{k}$:
@@
$\approx\sum_{\pi\in\Pi_{k}}\prod_{j=1}^{k}P(i_{\pi(j)}\mid b,h_{j-1}^{(\pi)},I)\cdot P(h_{j}^{(\pi)}\mid b,i_{\pi(j)},h_{j-1}^{(\pi)}),$ (22)
where $h_{0}^{(\pi)}=\emptyset$, and $k>=1$. Here, $h_{j}^{(\pi)}$ denotes the intermediate hypothesis state at step $j$ in the permutation $\pi$, which results from sequentially incorporating inspirations in the order $\{i_{\pi(1)},\ldots,i_{\pi(j)}\}$.
A.2 MOOSE-Chem Overview I/O Figure
Figure 2 shows the input and output overview of the MOOSE-Chem framework.

### v6

@@ Abstract
Scientific discovery plays a pivotal role in advancing human society, and recent progress in large language models (LLMs) suggests their potential to accelerate this process. However, it remains unclear whether LLMs can autonomously generate novel and valid hypotheses in chemistry. In this work, we investigate whether LLMs can discover high-quality chemistry hypotheses given only a research background—comprising a question and/or a survey—without restriction on the domain of the question. We begin with the observation that hypothesis discovery is a seemingly intractable task. To address this, we propose a formal mathematical decomposition grounded in a fundamental assumption: that most chemistry hypotheses can be composed from a research background and a set of inspirations. This decomposition leads to three practical subtasks—retrieving inspirations, composing hypotheses with inspirations, and ranking hypotheses—which together constitute a sufficient set of subtasks for the overall scientific discovery task. We further develop an agentic LLM framework, MOOSE-Chem, that is a direct implementation of this mathematical decomposition. To evaluate this framework, we construct a benchmark of 51 high-impact chemistry papers published and online after January 2024, each manually annotated by PhD chemists with background, inspirations, and hypothesis. The framework is able to rediscover many hypotheses with high similarity to the groundtruth, successfully capturing the core innovations—while ensuring no data contamination since it uses an LLM with knowledge cutoff date prior to 2024. Finally, based on LLM’s surprisingly high accuracy on inspiration retrieval, a task with inherently out-of-distribution nature, we propose a bold assumption: that LLMs may already encode latent scientific knowledge associations not yet recognized by humans.  [^1] All code and data can be found in https://github.com/ZonglinY/MOOSE-ChemContribution during internship at Shanghai Artificial Intelligence Laboratory. Corresponding author.
@@ 1 Introduction
We design experiments with the benchmark to test the three fundamental questions and find that LLMs are highly capable. We also test MOOSE-Chem with the benchmark, mimicking the setting to run it in the wild by only giving a background and a corpus of up to 3,000 chemistry papers to select inspiration. Even in this challenging setting, MOOSE-Chem can still rediscover many hypotheses with very high similarity with the ground truth ones, covering the main innovations.
@@
[…]
@@ 3 Benchmark Construction
To achieve the goals, we split each collected paper into the following components: <<background question, background question (strict), background survey, background survey (strict), one to three inspiration paper titles and their reason to serve as an inspiration, research hypothesis, experiments, reasoning process, summarization of inspirations>>. Every component is described by text.
@@ 4.1 Fundamental Assumption and Following Decomposition
$P(h\mid b)\approx\prod_{j=1}^{k}P(i_{j}\mid b,h_{j-1},I)\cdot P(h_{j}\mid b,h_{j-1},i_{j}),\text{ where }h_{0}=\emptyset$ (2)
This decomposition naturally defines a Markov Decision Process (MDP), where each intermediate hypothesis $h_{j-1}$ and selected inspiration $i_{j}$ constitute a state–action pair, $P(i_{j}\mid b,h_{j-1},I)$ models the policy of selecting the next inspiration, and $P(h_{j}\mid b,h_{j-1},i_{j})$ models the state transition toward the next step of hypothesis. Here, $I$ denotes the full (chemistry) literature, representing the entire inspiration space to search for each $i$. The complete derivation and theoretical justification are provided in § A.2, which forms the core of this paper.
Equation 2 is meaningful in that by decomposing $P(h\mid b)$ into more practical and executable smaller questions, the seemingly impossible-to-solve $P(h\mid b)$ itself becomes practical. We analyze how $P(i_{j}\mid b,h_{j-1},I)$ and $P(h_{j}\mid b,h_{j-1},i_{j})$ are practical and executable by LLMs in § 5.1 and § 5.2 correspondingly.
@@
Supported by Equation 2 and Equation 3, as a result, to model $P(h\mid b)$, the only three components we need to model are $P(i_{j}\mid b,h_{j-1},I)$, $P(h_{j}\mid b,h_{j-1},i_{j})$, and $R(h)$. The implementation details of the three components are illustrated in the remaining subsections in § 4. Analyses of LLM’s ability on the three components are provided in § 5.
@@ 4.2.1 The General Picture
Our methodology is developed based on the fundamental assumption discussed in § 4.1. Specifically, we use LLMs to perform $P(i_{j}\mid b,h_{j-1},I)$, $P(h_{j}\mid b,h_{j-1},i_{j})$, and $R(h)$, and organize them into a multi-agent LLM-based framework. The input to the framework is only a background question and/or background survey, together with a (large) chemistry literature corpus to search for inspiration. The output of the framework is a list of ranked research hypothesis.
@@ 4.2.2 Design Details of $P(i_{j}\mid b,h_{j-1},I)$ And Its Motivation
4.2.3 Design Details of $P(h_{j}\mid b,h_{j-1},i_{j})$ And Its Motivation
@@ 5 Investigation on Fundamental Questions
$P(h\mid b)$ can be understood as the task to discover high-quality chemistry research hypothesis, given only a background question and/or background survey. Our central question to investigate is how well LLMs can perform $P(h\mid b)$. Supported by Equation 2 and 3, we break up this main question into three smaller questions: how well can LLMs perform (1) $P(i_{j}\mid b,h_{j-1},I)$, (2) $P(h_{j}\mid b,h_{j-1},i_{j})$, and (3) $R(h)$? All experiments are performed by GPT-4o (its training data is up to October 2023).
@@ 5.1 How Well can LLMs perform $P(i_{j}\mid b,h_{j-1},I)$?
5.2 How Well can LLMs perform $P(h_{j}\mid b,h_{j-1},i_{j})$?
5 points
Generated hypothesis covers three key points (or covers all the key points) and leverage
them similarly as in the groundtruth hypothesis; Extra key points do not have apparent flaws.
4 points
Generated hypothesis covers three key points (or covers all the key points) and leverage
them similarly as in the groundtruth hypothesis; Extra key points have apparent flaws.
3 points
Generated hypothesis covers two key points and leverage them similarly
as in the groundtruth hypothesis, but does not cover more or all key points
2 points
Generated hypothesis covers one key point and leverage it similarly
as in the groundtruth hypothesis, but does not cover more or all key points
1 point
Generated hypothesis covers at least one key point, but is used differently
as in the groundtruth hypothesis
0 point
Generated hypothesis does not cover any key point
@@ 6.1 Baselines
is a hypothesis discovery framework for the NLP and biochemical domain. It relies on semantic and citation neighbors to retrieve information to assist the background. As a result, the retrieved information could be very related to the background that might not be able to serve as an inspiration. Here, we implement SciMON with LLM-based inspiration retrieval.
@@ 6.2 Results
Table 11: MOOSE-Chem runs with |I||I|=300, mimicking the copilot setting. This table shows the statistics of the top Matched Score across the benchmark. The evaluation is done by experts.
@@ 7 Conclusion
We investigated this central question: “Can LLMs automatically discover novel and valid chemistry (including material science) research hypotheses (even those which deserve a publication in Nature, Science, or a similar level) given only a chemistry research background (consisting of a research question and/or a background survey), without limitation on the domain of the research question?”. We proposed a fundamental assumption to break up this seemingly impossible-to-solve central question into three smaller, more practical, and executable fundamental questions. Then, we investigated LLM’s ability on each of them. To this end, we constructed a benchmark consisting of chemistry and material science papers published and only be public in 2024. We also developed an LLM-based multi-agent framework consisting of three stages reflecting the three smaller fundamental questions. Experiments showed that the framework (runs in a copilot in-the-wild setting, with LLMs with training data up to October 2023) can rediscover many hypotheses with very high similarity with the ground-truth ones, covering the main innovations.
@@ Appendix A Appendix
A.1 MOOSE-Chem Overview I/O Figure
Figure 2 presents an overview of the input and output of the MOOSE-Chem framework. In this work, the inspiration corpus is initialized using the titles and abstracts of numerous papers, which together constitute the search space.
A.2 Proof of the Rigorous Decomposition of $P(h\mid b)$ Based on the Fundamental Assumption
@@
This assumption is reminiscent of Swanson Linking (Swanson, 1986) in the domain of literature-based discovery (LBD), also known as the “ABC model”, where two concepts A and C are hypothesized as linked if they both co-occur with some intermediate concept B in papers. Our assumption differs in that: (1) for a chemistry hypothesis published in a good venue, usually more than one inspiration is needed; (2) background and inspiration are not necessarily linked by a path of intermediate papers; (3) our assumption is applied to a majority of existing published chemistry hypotheses, while LBD has been considered to only focus on a very specific, narrow type of hypothesis (Wang et al., 2024b). It might indicate that a similar proportion of future chemistry hypotheses can also result from linkages of existing literature.
@@
Here, $k\in\mathbb{Z}$ represents the number of inspirations needed for a particular $h$. Typically in chemistry, $k\in[1,3]$. We further assume each hypothesis $h$ has a unique minimal set of inspirations $\{i_{1},\ldots,i_{k}\}$ that, combined with the background $b$, determines its formation. We refer to this as the uniqueness assumption.
@@
$P(h\mid b)$ $=\sum_{i_{1},\ldots,i_{k}}P(h,i_{1},\ldots,i_{k}\mid b)$ (5)
$=\sum_{\pi\in\Pi_{k}}P\big(h,i_{\pi(1)},\ldots,i_{\pi(k)}\mid b\big)$ (6)
$=P(h,i_{1},\ldots,i_{k}\mid b)$ (7)
$=\begin{cases}P(h\mid b,i_{1})\cdot P(i_{1}\mid b,I)&\text{if }k=1\\ P(h\mid b,i_{1},\ldots,i_{k})\cdot\prod_{j=2}^{k}P(i_{j}\mid b,i_{1},\ldots,i_{j-1},I)\cdot P(i_{1}\mid b,I)&\text{if }k>1\end{cases}$ (8)
Equation 5 expands $P(h\mid b)$ by marginalizing over all possible inspiration sequences $(i_{1},\ldots,i_{k})$ that could, together with the background $b$, yield $h$. Equation 6 applies the uniqueness assumption—each hypothesis $h$ corresponds to a unique minimal set of inspirations $\{i_{1},\ldots,i_{k}\}$, leaving only the order of incorporation as a source of variability. Let $\Pi_{k}$ denote all permutations of $\{1,\ldots,k\}$; thus, the marginalization becomes a sum over $\pi\in\Pi_{k}$. Equation 7 follows from the fixed-order assumption (introduced here to simplify the equation), which posits a canonical constructive order of inspirations (i.e., $|\Pi_{k}|=1$), collapsing the sum to a single term. Finally, Equation 8 expands this joint distribution using the chain rule.
II denotes the full inspiration space—that is, the set of all possible knowledge pieces that could serve as an inspiration for generating a new hypothesis. This space includes not only all existing chemistry knowledge but also potentially relevant knowledge from other disciplines that could be leveraged in formulating novel chemistry hypotheses. However, computing over the full space $I$ is computationally infeasible. To make the problem tractable, we approximate $I$ with a large but manageable subset $\hat{I}$, consisting of approximately 3,000 top cited chemistry papers from the existing chemistry literature.
@@
$P(h_{k}\mid b,i_{1},\ldots,i_{k})$ $\approx P(h_{k}\mid b,f(b,i_{1},\ldots,i_{k-1}),i_{k})$ $\text{if }k>1$ (9)
$=P(h_{k}\mid b,h_{k-1},i_{k})$ $\text{if }k>1$ (10)
@@
$P(i_{k+1}\mid b,i_{1},\ldots,i_{k},I)$ $\approx P(i_{k+1}\mid b,f(b,i_{1},\ldots,i_{k}),I)$ $\text{if }k>1$ (11)
$=P(i_{k+1}\mid b,h_{k},I)$ $\text{if }k>1$ (12)
As a result, to achieve the final $h_{k}$, we need to obtain $\{h_{1},\ldots,h_{k-1}\}$ first (if $k>1$). In addition, viewing $h$ as a “state” and $i$ as an “action”, obtaining $h$ and $i$ through $P(h_{k}\mid b,h_{k-1},i_{k})$ and $P(i_{k+1}\mid b,h_{k},I)$ correspondingly indicates a Markov property: (1) a new $h$ depends only on $b$, its previous $h$, and the current $i$; and (2) an $i$ depends only on $b$, II, and the previous $h$.
Markov Decision Process (MDP) Formulation.
Building upon this Markov property, we now interpret the sequential formation of hypotheses as a decision process. We therefore formalize this process as a Markov Decision Process (MDP), where:
  • the state is $(b,h_{j-1})$, encoding the research background and the current intermediate hypothesis;
  • the action is the selected inspiration $i_{j}$;
  • the transition function is $P(h_{j}\mid b,h_{j-1},i_{j})$, yielding the next hypothesis state; and
  • the reward can be defined as the quality or plausibility of the resulting hypothesis $h_{j}$ (e.g., as judged by an LLM or domain expert).
@@
Each transition remains conditioned on the background knowledge $b$, though $b$ is omitted from the notation to emphasize the Markov structure of the progression. This MDP view implies that hypothesis discovery can be modeled as a sequential policy choosing inspirations and applying transition dynamics to update the intermediate hypothesis state. We next make this view explicit in the probability space.
Building on this formulation, we interpret the formation of a hypothesis $h$ (specifically, $h=h_{k}$) as a constructive process that sequentially integrates a set of inspirations $\{i_{1},\ldots,i_{k}\}$ into intermediate hypothesis states $\{h_{1},\ldots,h_{k}\}$.
Formally, the conditional probability $P(h\mid b)$ can be expressed as a marginal over all valid sequences of inspirations that can generate $h$ (supported by assumption in Equation 4 and uniqueness assumption):
@@
where $\Pi_{k}$ denotes the set of all permutations of $\{1,\ldots,k\}$ applied to the inspirations $\{i_{1},\ldots,i_{k}\}$ such that the resulting composition yields the final hypothesis $h_{k}=h$, under the assumptions of Equation 4 and the uniqueness assumption, which specify that $h$ is fully determined by $b$ and $\{i_{1},\ldots,i_{k}\}$. In this context, $h_{j}$ represents the intermediate hypothesis state obtained after integrating $i_{\pi(j)}$ at step $j$.
The degree to which the order of inspirations $\{i_{1},\ldots,i_{k}\}$ affects hypothesis formulation can vary across disciplines. In empirical sciences such as chemistry, the contributions of individual inspirations are largely interchangeable, and their order of integration has limited impact on the final hypothesis, resulting in a large $|\Pi_{k}|$. Conversely, in disciplines such as mathematics, where hypotheses (e.g., theorems) often require constructing a specific sequence of lemmas and prior results, the ordering of inspirations is more constrained and may follow a near-deterministic path.
For simplicity of exposition, we adopt the fixed-order assumption introduced earlier—i.e., $|\Pi_{k}|=1$—which selects a canonical constructive order of inspirations $\{i_{1},\ldots,i_{k}\}$ for analysis. Under this assumption, the hypothesis $h$ is constructed through that specific sequence, giving:
@@
Therefore, under the uniqueness assumption and fixed-order assumption, for $k>1$,
@@
$=\prod_{j=1}^{k}P(i_{j}\mid b,h_{j-1},I)\cdot P(h_{j}\mid b,h_{j-1},i_{j}),\text{ where }h_{0}=\emptyset$ (18)
Equation 15 follows from the assumption in Equation 4, together with the uniqueness and fixed-order assumptions, which specify that $h$ is determined by $b$ and a unique ordered set of inspirations $\{i_{1},\ldots,i_{k}\}$. Equation 16 applies the chain rule to factorize the joint distribution. Equation 17 follows from the Markov property, assuming that the next $(i_{j},h_{j})$ pair depends only on $b$ and the preceding state $h_{j-1}$. Finally, Equation 18 re-applies the chain rule, with $P(I)=1$ by definition.
Although starting from $k>1$, Derivation 18 covers the situation when $k=1$ in Equation 8. Therefore, in sum, we successfully break up the seemingly impossible question $P(h\mid b)$ into many practical and executable smaller questions as:
$P(h\mid b)\approx\prod_{j=1}^{k}P(i_{j}\mid b,h_{j-1},I)\cdot P(h_{j}\mid b,h_{j-1},i_{j}),\text{ where }h_{0}=\emptyset\text{ and }k\geq 1$ (19)
Equation 19 provides the MDP-based decomposition of $P(h\mid b)$, where each term $P(i_{j}\mid b,h_{j-1},I)$ and $P(h_{j}\mid b,h_{j-1},i_{j})$ corresponds respectively to the policy (action selection) and state transition components of the underlying Markov Decision Process. This formulation highlights that scientific hypothesis discovery can be viewed as a sequential decision-making problem, where an agent iteratively selects inspirations to maximize the expected quality (usually compounded evaluation of validity and novelty) of the final hypothesis $h_{k}$.
Of course, without the fixed-order assumption, a more complete derivation of $P(h\mid b)$ involves marginalizing over all valid permutations in $\Pi_{k}$:
@@
$\approx\sum_{\pi\in\Pi_{k}}\prod_{j=1}^{k}P(i_{\pi(j)}\mid b,h_{j-1}^{(\pi)},I)\cdot P(h_{j}^{(\pi)}\mid b,h_{j-1}^{(\pi)},i_{\pi(j)}),$ (21)
where $h_{0}^{(\pi)}=\emptyset$, $h_{k}^{(\pi)}=h$, and $k\geq 1$. Here, $h_{j}^{(\pi)}$ denotes the intermediate hypothesis state at step $j$ in the permutation $\pi$, which results from sequentially incorporating inspirations in the order $\{i_{\pi(1)},\ldots,i_{\pi(j)}\}$.